Solve

"\frac{dy}{dx}" =(x+2y-3)/(2x+y-3)




1
Expert's answer
2021-08-11T12:51:17-0400

Our standard toolkit for DE's cannot be used. However, we can perform a transformation to remove the constants from the linear numerator and denominator.

Consider the simultaneous equations

"\begin{cases}
 x+2y-3=0 \\
 2x+y-3=0
\end{cases} \implies \begin{cases}
 x=1\\
 y=1
\end{cases}"

As a result we perform two linear transformations:

"\begin{cases}
 u=x-1 \\
 v=y-1
\end{cases} \implies \begin{cases}
 x=u+1\\
 y=v+1
\end{cases} \implies \begin{cases}
 \frac{dx}{du}=1\\
 \frac{dy}{dv}=1
\end{cases}"

And if we substitute into the DE we get

"\frac{dy}{dx}=\frac{(u+1)+2(v+1)-3}{2(u+1)+2(v+1)-3}\\

\frac{dy}{dx}=\frac{u+1+2v+2-3}{2u+2+v+1-3}\\
\frac{dy}{dx}=\frac{u+2v}{2u+v}"

And utilising the chain rule we have

"\frac{dy}{dx}= \frac{dy}{dv}\frac{dv}{du} \frac{du}{dx} \implies \frac{dy}{dx}= \frac{dv}{du}"

Thus we have a transformed equation

"\frac{dv}{du}= \frac{u+2v}{2u+v}\\"


This is now in a form that we can handle using a substitution of form v =wu, which if we differentiate wrt u using the product gives us.

Using this substitution into our modified DE, we get


"u \frac{dw}{du}= \frac{1+w^2}{2+w}\\"

This is now a separable DE, so we can rearrange and separate the variables to get:

"\int \frac{1}{u}du = \int \frac{2+w}{1+w^2}\\
\ln
|
w
+
1
|
−
\ln
|
w
−
1
|
−
1
2
\ln
∣
∣
w
2
−
1
∣
∣
=
\ln
|
u
|
+
\ln
C"

"\frac{|
w
+
1
|}{
|
w
−
1
|
√
w
2
−
1}
=
|
C
u
|\\
\therefore (w+1)=Au^2(w-1)^3"

Then restoring the earlier w substitution, using "w

=

\frac{v}{



u}"  we get:

"\u2234\nv\n+\nu\n=\nA\n(\nv\n\u2212\nu\n)^\n3"

"(\ny\n\u2212\n1\n)\n+\n(\nx\n\u2212\n1\n)\n=\nA\n(\n(\ny\n\u2212\n1\n)\n\u2212\n(\nx\n\u2212\n1\n)\n)^\n3"

"\u2234\ny\n+\nx\n\u2212\n2\n=\nA\n(\ny\n\u2212\nx\n)^\n3"

This is the General Solution, in implicit form.